2 to the power of 6 = 4 to the power of n?
Answer by stanbon(48516)

2 to the power of 6 = 4 to the power of n?
----
2^6 = 4^n
---
2^6 = (2^2)^n
---
2^6 = 2^(2n)
---
2n = 6
n = 3
==============
Cheers,
Stan H.